How many grams of NaHCO3 would have to be added to 2.00 L of 0.100 M H2CO3 to yield a solution with a pH = 6.00? (Ka = 4.2x10-7

1
Expert's answer
2021-02-02T01:22:15-0500

Solution:

"pH=pK_a+log\dfrac{[HCO_2^-]}{[H_2CO_3]}"


"6=6.38+log\dfrac{[HCO_3^-]}{0.1}"


"[HCO_3^-]=0.042\;M"


"m=0.042\;M*2L*84.01(\frac{g}{mole})=7.06\;g\;of\;NaHCO_3"


Answer:


7.06 grams of NaHCO3 should be added.